OpenStudy (anonymous):

(x+1)^2 = x^1+2 + 1^1+2 = x^3 + 1^3 =x^3 + 1 Does this make sense? or am I missing something?

OpenStudy (anonymous):

na this does not make any sense (x+1)^2 means (x+1)*(x+1)

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

Okay, guess I'll try the other method then..ty

OpenStudy (anonymous):

i.e. x*(x+1) +1 *(x+1) = x*x+x*1+1*x+1*1 =x^2 +2x+1

OpenStudy (anonymous):

ah yep, got that :3 ty Vedic